If one angle of a pentagon is a right angle. find the size of each of the remaining angles if they are equal to one another.
Answer by ikleyn(52847)

.
Hint. The sum of interior angles of any pentagon is (5-2)*180 degs = 3*180 degs = 540 degs.

To get the final answer, subtract 90 degs and divide the difference by 4.